OpenStudy (anonymous):

choose the correct simplification of (x^3/y^5)^2

OpenStudy (kamibug):

What are the options?

OpenStudy (anonymous):

a,)x^5/y^7 b.)x/y^3 c.)x^6/y^10 d,)x^6y^10

OpenStudy (agent0smith):

Try it yourself first. \[\Large (x^a)^b = x^{ab}\]

OpenStudy (anonymous):

you add them?

OpenStudy (anonymous):

you add 3 and 2 then you add 5 and 2?

OpenStudy (anonymous):

or do you multiply them? 3.2 and 5.2

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{x^3}{y^5}\right)^{\color{blue}{ 2}}\implies \cfrac{x^{3\cdot {\color{blue}{ 2}}}}{y^{5\cdot {\color{blue}{ 2}}}}\)

OpenStudy (agent0smith):

\[\Large (x^a)^b = x^{ab}\]ab means a multiplied by b. Addition would require an... addition sign.

OpenStudy (anonymous):

so the answer would be x^6 over y^10?

OpenStudy (anonymous):

@agent0smith @jdoe0001

OpenStudy (jdoe0001):

so the answer would be x^6 over y^10? \(\Large \checkmark\)

OpenStudy (anonymous):

choose the correct simplification of 8x^2(5x+2x^2-3)

OpenStudy (jdoe0001):

just distribute I'd think